# Apparent and emergent dark matter around a Schwarzschild black hole

### v1

@@ I Introduction
Through the vast expanse of the cosmos, two enigmatic entities dominate our understanding of the universe: black holes (BH) and dark matter (DM). Black holes, the remnants of massive stars that have collapsed under their own gravity, continue to captivate the scientific community and the public imagination with their gravitational prowess, capable of bending space and time beyond comprehension. Despite their theoretical prediction by Einstein’s general theory of relativity Schwarzschild (1916), the direct evidence for black holes remained elusive for decades until the pioneering work of astronomers like Andrea Ghez and Reinhard Genzel, who observed stars orbiting an invisible object at the center of our Milky Way, providing compelling evidence for the existence of supermassive black holes Ghez et al. (2008). On April 10, 2019, the Event Horizon Telescope (EHT) collaboration obtained the first electromagnetic and polarized-based image of the black hole shadow at the heart of Messier 87 Akiyama et al. (2019a, b, 2022a). Then, the image of the supermassive black hole (SMBH) black hole shadow at the heart of our galaxy, the Milky Way, followed Akiyama et al. (2022b, c), confirming the existence of these enigmatic objects.
@@
In parallel, the cosmic census has revealed an astonishing truth: the visible matter that we are made of - stars, planets, galaxies - comprises only a fraction of the total matter content of the universe. The majority, approximately 85% de Vries (1999), is concealed within the shroud of darkness known as dark matter. Yet, despite its pervasive presence, dark matter remains elusive, evading detection through conventional means due to its lack of interaction with electromagnetic radiation. Various theoretical models have been proposed in the quest to unravel the enigma of dark matter, each offering unique insights into its composition and properties (For reviews, see Refs. Garrett and Duda (2011); Ureña López (2019); Arun et al. (2017); Arbey and Mahmoudi (2021)). Associated with the $\Lambda$CDM model is the detection of the Weakly Interacting Massive Particles (WIMP) or axions de Swart et al. (2017). Scientists have devised ingenious detection methods to probe the mysteries of dark matter, ranging from direct searches using underground detectors Baum et al. (2020) to indirect approaches through astronomical observations and particle accelerators. However, despite decades of dedicated efforts that sometimes seemingly show positive results Bernabei et al. (2008, 2013, 2018), the elusive nature of dark matter persists, debunking claims for detecting it Angloher et al. (2016); Amole et al. (2017); Aalbers et al. (2022), and continues to challenge our fundamental understanding of the cosmos. Not only these, but the model is plagued with certain anomalies at the astrophysical level Moore (1994); Klypin et al. (1999); Boylan-Kolchin et al. (2011).
@@
Recent advancements in astrophysical instrumentation, theoretical frameworks, and phenomenologies have unveiled a tantalizing possibility: an SMBH acting as a laboratory to detect unique imprints of dark matter that surrounds it. In the literature, several researchers have attempted to go in this direction Xu et al. (2018); Hou et al. (2018a, b); Konoplya (2019); Haroon et al. (2019); Xu et al. (2020); Jusufi et al. (2020); Xu et al. (2021); Nampalliwar et al. (2021); Jusufi and Saurabh (2021); Konoplya (2021); Saurabh and Jusufi (2021); Pantig and Övgün (2022a, b); Atamurotov et al. (2022); Jusufi (2023); Konoplya and Zhidenko (2022); Pantig et al. (2022); Liu et al. (2022); Pantig and Övgün (2022c); Anjum et al. (2023); Övgün et al. (2023); Errehymy et al. (2023); Qiao and Zhou (2023); Zhou et al. (2023); Capozziello et al. (2023a, b); Liu et al. (2023); Della Monica and de Martino (2023); Yang et al. (2024); Gómez and Valageas (2024); Qiao and Su (2024); Wu et al. (2024); Macedo et al. (2024). Indeed, when more sophisticated astrophysical instrumentation becomes available in the future, black holes may potentially unveil the true nature of dark matter.
@@ IV Null geodesic, black hole shadow and EHT constraints
First, we examine plots derived from theoretical considerations, shown in the left panel of Fig. 3. In these plots, the observer-dependence of the shadow radius is also evident. Generally, the shadow radius is small when the observer is near the black hole. The upper left panel shows $R_{\rm sh}$ curve under the influence of the non-locality of emergent gravity. An observer around the range $r_{\rm o}$ may perceive a larger shadow for $\xi=0.1$ than $\xi=0.3$. As the observer goes very far, the effect reverses. Higher values of $\xi$ make the observer perceive the Schwarzschild shadow at a closer distance from the black hole. Interestingly, while the photonsphere radius decreases while $\xi$ increases, the reverse effect happens to the shadow radius. The parameter $\xi$ has been shown to amplify the difference between the shadow radius from $\Delta\xi=0.1$.
The emergent dark matter arising from dark energy modeled as a superconductor has shown a different behavior for the shadow radius. See the lower left panel. First, both $r_{\rm ps}$ and $R_{\rm sh}$ increases with $\eta$. Comparing this to the apparent dark matter from the non-local emergent gravity model, $\eta$ does not show any amplification between two values of $R_{\rm sh}$ even when the observer distance from the black hole is large, which is a different behavior than the one we observe for the photonsphere. Interestingly enough, for a given value of $\eta$, increasing the value of the screening length parameter increases $r_{\rm ps}$ but is shown to decrease $R_{\rm sh}$.
@@
Figure 3: The figures on the left panel show the theoretical behavior of the shadow radius that is observer-dependent. The vertical red and horizontal black dotted lines correspond to the Schwarzschild horizon and shadow. The right panel shows the constraints on $\xi$ and $\eta$ using actual data from EHT for Sgr. A* (blue shade) and M87* (red shade).
A final caveat: while actual black holes have spin parameters described by the parameter $a$, constraining $\xi$ and $\eta$ using the non-rotating solution might give unrealistic results. However, it was explained in Ref. Vagnozzi et al. (2023) why considering the non-rotating case is sufficient in finding constraints for parameters affecting the black hole geometry. Furthermore, we also neglect photons escaping from the horizon itself. Hence, the calculations above only consider the silhouette of the invisible black hole shadow (see Ref. Dokuchaev and Nazarova (2020)), which is prevalent in the literature.
@@ V Conclusion
Intrigued by how these two dark matter models affect the black hole geometry at the center of a galaxy, we investigate their effects on the event horizon, null geodesic, and the invisible black hole shadow. We found unique descriptions and differences between the two models, as shown by the plots in Figs. 1-3, which we discussed in Sect. IV. Perhaps one of the intriguing results from this study is that the effect of the parameter $l$ can be considered negligible when applied to SMBHs. Meanwhile, it was shown how this length scale differs from the screening length parameter $\lambda_{\rm G}$ as its effect is more pronounced.
We also found upper bounds to these parameters, which are already depicted in the right panel of Fig. 3. It is a useful plot since it allows us to estimate variables related to $\xi$ and $\eta$. For instance, since is defined as $\xi^{2}=(a_{0}/6)GM$, it has a dimension of $[\text{L}^{2}]/[\text{s}^{2}]$. Thus, if we have an upper bound for M87* as $\xi=0.019$, this is equivalent to $\xi=1.708\times 10^{15}\text{ m}^{2}/\text{s}^{2}$. The corresponding value for $M$ is then $\sim 4.91\times 10^{50}\text{ kg}\sim 2.47\times 10^{20}M_{\odot}$ at $3\sigma$ level. We could also do the same for $\eta$ but $M$ will depend on what specific value for $\alpha$ there is (or vice-versa). We may have lower estimates for $M$ at lower $\sigma$ levels. For instance, in Refs. Merritt and Tremblay (1993); Murphy et al. (2011), the estimated total mass within the halo ($40-50$ kpc) is around $\sim 6\times 10^{12}M_{\odot}$ at 90% confidence.

### v2

@@ I Introduction
Through the vast expanse of the cosmos, two enigmatic entities dominate our understanding of the universe: black holes (BH) and dark matter (DM). Black holes, the remnants of massive stars that have collapsed under their own gravity, continue to captivate the scientific community and the public imagination with their gravitational prowess, capable of bending space and time beyond comprehension. Despite their theoretical prediction by Einstein’s general theory of relativity Schwarzschild (1916), the direct evidence for black holes remained elusive for decades until the pioneering work of astronomers like Andrea Ghez and Reinhard Genzel, who observed stars orbiting an invisible object at the center of our Milky Way, providing compelling evidence for the existence of supermassive black holes Ghez et al. (2008). On April 10, 2019, the Event Horizon Telescope (EHT) collaboration obtained the first electromagnetic and polarized-based image of the black hole shadow at the heart of Messier 87 Akiyama et al. (2019a, b, 2022a). Then, the image of the supermassive black hole (SMBH) black hole shadow at the heart of our galaxy, the Milky Way, followed Akiyama et al. (2022b, c), confirming the existence of these enigmatic objects. Not only that these results confirm the seminal works by Synge, Luminet, and Falcke Synge (1966); Luminet (1979); Falcke et al. (2000), but they also sparked great interest among the scientific community to study the black hole shadow in various black hole models coming from different phenomenologies and alternative theories of gravity (See Refs. Bambi et al. (2019); Zahid et al. (2022); Rayimbaev et al. (2022); Mustafa et al. (2022a); Afrin et al. (2023); Zubair et al. (2023); Zahid et al. (2023); Ghorani et al. (2023); Sekhmani et al. (2024a); Raza et al. (2024) to refer a few).
@@
In parallel, the cosmic census has revealed an astonishing truth: the visible matter that we are made of - stars, planets, galaxies - comprises only a fraction of the total matter content of the universe. The majority, approximately 85% de Vries (1999), is concealed within the shroud of darkness known as dark matter. Yet, despite its pervasive presence, dark matter remains elusive, evading detection through conventional means due to its lack of interaction with electromagnetic radiation. Various theoretical models have been proposed in the quest to unravel the enigma of dark matter, each offering unique insights into its composition and properties (For an excellent and comprehensive reviews, see Refs. Garrett and Duda (2011); Ureña López (2019); Arun et al. (2017); Arbey and Mahmoudi (2021)). Associated with the $\Lambda$CDM model is the detection of the Weakly Interacting Massive Particles (WIMP) or axions de Swart et al. (2017). Scientists have devised ingenious detection methods to probe the mysteries of dark matter, ranging from direct searches using underground detectors Baum et al. (2020) to indirect approaches through astronomical observations and particle accelerators. However, despite decades of dedicated efforts that sometimes seemingly show positive results Bernabei et al. (2008, 2013, 2018), the elusive nature of dark matter persists, debunking claims for detecting it Angloher et al. (2016); Amole et al. (2017); Aalbers et al. (2022), and continues to challenge our fundamental understanding of the cosmos. Not only these, but the model is plagued with certain anomalies at the astrophysical level Moore (1994); Klypin et al. (1999); Boylan-Kolchin et al. (2011).
@@
Recent advancements in astrophysical instrumentation, theoretical frameworks, and phenomenologies have unveiled a tantalizing possibility: an SMBH acting as a laboratory to detect unique imprints of dark matter that surrounds it. In the literature, several researchers have attempted to go in this direction, particularly studying DM effects on the black hole shadow Hou et al. (2018a, b); Konoplya (2019); Jusufi et al. (2020); Nampalliwar et al. (2021); Pantig and Övgün (2022a); Jusufi and Saurabh (2021); Jusufi (2023); Pantig and Övgün (2022b); Anjum et al. (2023); Övgün et al. (2023); Capozziello et al. (2023a, b); Yang et al. (2024); Gómez and Valageas (2024); Wu et al. (2024), gravitational lensing effects Haroon et al. (2019); Pantig and Övgün (2022c); Atamurotov et al. (2022); Pantig et al. (2022); Qiao and Zhou (2023); Liu et al. (2023); Macedo et al. (2024), null structures such as event horizon and ergosphere Xu et al. (2018, 2020, 2021), quasinormal modes and greybody factors Konoplya (2021), thermodynamic properties Konoplya (2021); Errehymy et al. (2023); Rakhimova et al. (2023); Sekhmani et al. (2024b), spherical accretions Saurabh and Jusufi (2021), super-radiant instabilities Liu et al. (2022), wormhole solutions and their analysis Mustafa et al. (2022b, 2024), and others Della Monica and de Martino (2023); Qiao and Su (2024); Konoplya and Zhidenko (2022); Zhou et al. (2023); Rayimbaev et al. (2021). Indeed, when more sophisticated astrophysical instrumentation becomes available in the future, black holes may potentially unveil the true nature of dark matter.
@@ IV Null geodesic, black hole shadow and EHT constraints
We examine plots derived from theoretical considerations, shown in the left panel of Fig. 3. In these plots, the observer-dependence of the shadow radius is also evident. Generally, the shadow radius is small when the observer is near the black hole. The upper left panel shows $R_{\rm sh}$ curve under the influence of the non-locality of emergent gravity. An observer around the range $r_{\rm o}$ may perceive a larger shadow for $\xi=0.1$ than $\xi=0.3$. As the observer goes very far from the black hole, the effect reverses. Higher values of $\xi$ make the observer perceive the Schwarzschild shadow at a closer distance from the black hole. Interestingly, while the photonsphere radius decreases while $\xi$ increases, the reverse effect happens to the shadow radius. The parameter $\xi$ has been shown to amplify the difference between the shadow radius from $\Delta\xi=0.1$. The shadow cast for this case is shown in the upper right panel, plotted once the numerical value for $R_{\rm sh}$ is found as we approximate $r_{\rm o}\rightarrow\infty$.
The emergent dark matter arising from dark energy modeled as a superconductor has shown a different behavior for the shadow radius. See the lower left panel. First, both $r_{\rm ps}$ and $R_{\rm sh}$ increases with $\eta$. Comparing this to the apparent dark matter from the non-local emergent gravity model, $\eta$ does not show any amplification between two values of $R_{\rm sh}$ even when the observer distance from the black hole is large, which is a different behavior than the one we observe for the photonsphere. Interestingly enough, for a given value of $\eta$, increasing the value of the screening length parameter increases $r_{\rm ps}$ but is shown to decrease $R_{\rm sh}$. The shadow cast plotted in the lower right panel agrees with the observation from the remote observer.
Figure 3: The figures on the left panel show the theoretical behavior of the shadow radius that is observer-dependent. The vertical red and horizontal black dotted lines correspond to the Schwarzschild horizon and shadow. The right panel shows the shadow cast as perceived by some remote observer.
@@
Figure 4: Constraints on $\xi$ and $\eta$ using actual data from EHT for Sgr. A* (blue shade) and M87* (red shade). The annotated values indicated are the upper bounds. The lower bound is always zero, representing the Schwarzschild case not surrounded by dark matter.
A final caveat: while actual black holes have spin parameters described by the parameter $a$, constraining $\xi$ and $\eta$ using the non-rotating solution might give unrealistic results. However, it was explained in Ref. Vagnozzi et al. (2023) why considering the non-rotating case is sufficient in finding constraints for parameters affecting the black hole geometry. Furthermore, we also neglect radiation escaping near the event horizon. Hence, the calculations above only consider the silhouette of the invisible black hole shadow (see Ref. Dokuchaev and Nazarova (2020)), which is prevalent in the literature.
@@ V Conclusion
Intrigued by how these two dark matter models affect the black hole geometry at the center of a galaxy, we investigate their effects on the event horizon, null geodesic, and the invisible black hole shadow. We found unique descriptions and differences between the two models, as shown by the plots in Figs. 1-4, which we discussed in Sect. IV. Perhaps one of the intriguing results from this study is that the effect of the parameter $l$ can be considered negligible when applied to SMBHs. Meanwhile, it was shown how this length scale differs from the screening length parameter $\lambda_{\rm G}$ as its effect is more pronounced.
We also found upper bounds to these parameters, which are already depicted in the right panel of Fig. 4. It is a useful plot since it allows us to estimate variables related to $\xi$ and $\eta$. For instance, since the latter is defined as $\xi^{2}=(a_{0}/6)GM$, it has a dimension of $[\text{L}]^{2}/[\text{t}]^{2}$. Thus, if we have an upper bound for M87* as $\xi=0.019$, multiplying by $c^{2}$ gives $\xi=1.708\times 10^{15}\text{ m}^{2}/\text{s}^{2}$. The corresponding value for DM mass $M$ is then $\sim 4.91\times 10^{50}\text{ kg}\sim 2.47\times 10^{20}M_{\odot}$ at $3\sigma$ level. We could also do the same for $\eta$, but $M$ will depend on the specific value for $\alpha$ (or vice-versa). We may have lower estimates for $M$ at lower $\sigma$ levels. In Refs. Merritt and Tremblay (1993); Murphy et al. (2011), the estimated total mass within the halo restricted from the range $40-50$ kpc is around $\sim 6\times 10^{12}M_{\odot}$ at 90% confidence. While the model considered in this paper does not specify the range of halo concentration, the higher mass $M$ requirement only indicates that a large concentration is needed to affect the shadow.